OpenStudy (anonymous):

If the midpoint of segment AB is 0,0 and if the coordinates of point A are (x,y) then coordinates of point B are

OpenStudy (anonymous):

(-x,-y ) (-x,y) (x,-y) (x,y) (2x, 2y)

OpenStudy (anonymous):

@jim_thompson5910 can you help?

OpenStudy (anonymous):

If M(0, 0) is the midpoint of two points A(x, y) and B(x1, y1), then 0 = (x + x1)/2 and 0 = (y + y1)/2.

OpenStudy (campbell_st):

perhaps you should draw a picture to help understand 

OpenStudy (anonymous):

So it's -x, -y?

OpenStudy (anonymous):

yeah eaxctly..... i guess!!! :P

OpenStudy (campbell_st):

that's correct,